The following results were obtained as part of a multiple regression analysis involving 3 independent variables:

SSR = 12240, SSYY = 16900,    and n = 31

          The calculated value of the F statistic is: (6 pts)
Answer #1
Sources Sum of Square Degree of freedom Mean square F value
Regression 12240 3-1 = 2 12240/2 = 6120 F = 6120/166.43 = 36.77
Error 16900 - 12240 = 4660 30-2 = 28 4660/28 = 166.43
Total 16900 31-1 = 30